A curve has an x-intercept at * = 1. At its y-intercept of y = -3, the curve's tangent line

has a slope of -1. Find the equation of this curve in the form y

= ax2 + hx+c


1
Expert's answer
2022-03-25T05:50:14-0400

"y=ax^2+bx+c\\\\y=0,x=1\\Rightarrow a+b+c=0\\\\x=0,y=-3\\Rightarrow c=-3\\\\y'\\left( 0 \\right) =b=-1\\\\\\left\\{ \\begin{array}{c}\ta+b+c=0\\\\\tc=-3\\\\\tb=-1\\\\\\end{array} \\right. \\Rightarrow a=4,b=-1,c=-3\\\\y=4x^2-x-3"